OpenStudy (anonymous):

Solve 5^x + 5 = 9^x anyone help please :(

OpenStudy (anonymous):

first u subtract 9x-5x then u should get 5=4x then u divide and u get 5/4=x

OpenStudy (anonymous):

x ≈ –3.63 x ≈ –0.28 x ≈ 0.07 x ≈ 13.69 those are my answers

OpenStudy (anonymous):

r u sure it has to be one of those?

OpenStudy (anonymous):

positive it is solving exponential equations with unequal bases

OpenStudy (anonymous):

omg im so sorry I wrote the question wrong it is 5^x+5=9^x

OpenStudy (anonymous):

oh now it makes sense lol. ok let me try again

OpenStudy (anonymous):

it is \[\large 5^{x+5}=9^x\] maybe ?

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

then start with \[(x+5)\ln(5)=x\ln(9)\] and do some algebra to solve for \(x\)

OpenStudy (anonymous):

don't be confused by the \(\ln(5)\) and \(\ln(9)\), they are just numbers

OpenStudy (anonymous):

im so confused could you help me out

OpenStudy (anonymous):

multiply out and get \[x\ln(5)+5\ln(5)=x\ln(9)\] subtract \(x\ln(5)\) and get \[5\ln(5)=x\ln(9)-5\ln(5)\] factor out the \(x\) for \[5\ln(5)=(\ln(9)-\ln(5))x\] and finally divide to get \[\frac{5\ln(5)}{\ln(9)-\ln(5)}=x\]

OpenStudy (anonymous):

i understand what you did but i don't understand the concept of ln

OpenStudy (anonymous):

so i wouldn't know how to divide that :(